# Cartesian Nodal Lines and Magnetic Kramers Weyl Nodes in Spin-Split Antiferromagnets

### v1

@@ Abstract
Here, $G_{z}$ represents the reciprocal lattice vector along the $z$ direction. If the lattice constant is restored, the expression of $\sigma_{xy}$ is ±$\pm\frac{e^{2}}{h}\frac{1}{a_{z}}$, where $a_{z}$ represents the lattice constant along the $z$ direction. This quantization requires that all $k_{z}$ planes except for $k_{z}=0$ and $k_{z}=\pi$ are gapped. In other words, it requires that the Fermi surface corresponds to the two nodal rings at the $k_{z}=0$ and $k_{z}=\pi$ planes. In our model, this corresponds to the conditions that $t=t_{z}=0$ and $\mu=0$ are to be satisfied.
Weyl node position Charge Requirements for existance
$\left(0,0,-\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ $+1$ $|B_{z}|<\lambda_{\rm so}$
$\left(0,0,-\pi+\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ $-1$
$\left(\pi,\pi,-\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ $+1$
$\left(\pi,\pi,-\pi+\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ $-1$
$\left(0,\pi,-\arcsin\frac{B_{z}+2\lambda_{\rm M}\eta}{\lambda_{\rm so}}\right)$ $-1$ $\eta<\eta_{c}-\frac{B_{z}}{2\lambda_{\rm M}}$
$\left(0,\pi,-\pi+\arcsin\frac{B_{z}+2\lambda_{\rm M}\eta}{\lambda_{\rm so}}\right)$ $+1$
$\left(\pi,0,\arcsin\frac{2\lambda_{\rm M}\eta-B_{z}}{\lambda_{\rm so}}\right)$ $-1$ $\eta<\eta_{c}+\frac{B_{z}}{2\lambda_{\rm M}}$
$\left(\pi,0,\pi-\arcsin\frac{2\lambda_{\rm M}\eta-B_{z}}{\lambda_{\rm so}}\right)$ $+1$
(±$\left(\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),0,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $-1$ $\lambda_{\rm so}<\lambda_{\rm M}$, -$\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}<B_{z}<\lambda_{\rm M}(2\eta-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})$
(±$\left(\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),0,\pi-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $+1$
(±$\left(\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),\pi,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $-1$ $\lambda_{\rm so}<\lambda_{\rm M}$, -$\lambda_{\rm M}(2\eta+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})<B_{z}<-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}$
(±$\left(\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),\pi,\pi-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $+1$
$\left(\pi,\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $-1$ $\lambda_{\rm so}<\lambda_{\rm M}$, $\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}<B_{z}<\lambda_{\rm M}(2\eta+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})$
$\left(\pi,\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),-\pi+\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $+1$
$\left(0,\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $-1$ $\lambda_{\rm so}<\lambda_{\rm M}$, $-\lambda_{\rm M}(2\eta-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})<B_{z}<\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}$
$\left(0,\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),-\pi+\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ $+1$
Table 1: Position, charge, and existence conditions for Weyl nodes under a Zeeman field in the $z$ direction. $\eta_{c}=\lambda_{\rm so}/2\lambda_{\rm M}$.
@@
Intriguingly, if $t$ and $t_{z}$ remain zero, we observe that an approximate Hall plateau can emerge even for finite $\mu$, provided that the condition $B_{z}>\mu$ is satisfied. This phenomenon arises because the energy gaps in the $k_{z}$ planes (for $k_{z}\neq 0$ and $k_{z}=\pi$) increase as $B_{z}$. When $\mu$ lies within these energy gaps, the contributions from the $k_{z}$ planes become quantized. However, since this system is a semimetal, the energy gaps of the $k_{z}$ planes varies continuously as a function of $k_{z}$. A larger $B_{z}$ enhances the $k_{z}$-dependence of this function, making it more sharply varying. Nevertheless, planes sufficiently close to $k_{z}=0$ and $k_{z}=\pi$ always exhibit nonzero and nonquantized contributions. Consequently, the plateau is approximate as long as $\mu$ becomes finite.
@@
In real materials, $t$ and $t_{z}$ are generally finite, making the observation of the predicted Hall plateau less realistic. Nevertheless, a strong anomalous Hall effect driven by a weak field remains an intriguing and experimentally observable phenomenon. As a final remark, since the direction of the Zeeman field strongly influences both symmetry and band structure, the anomalous Hall effect will exhibit a specific angular dependence. This angular dependence can serve as an additional method to diagnose the band structure, complementing techniques such as angle-resolved photoemission spectroscopy.

### v3

@@ Abstract
Here, $G_{z}$ represents the reciprocal lattice vector along the $z$ direction. If the lattice constant is restored, the expression of $\sigma_{xy}$ is $\pm\frac{e^{2}}{h}\frac{1}{a_{z}}$, where $a_{z}$ represents the lattice constant along the $z$ direction. This quantization requires that all $k_{z}$ planes except for $k_{z}=0$ and $k_{z}=\pi$ are gapped. In other words, it requires that the Fermi surface coincides with the two nodal rings at the $k_{z}=0$ and $k_{z}=\pi$ planes. In our model, this corresponds to the conditions that $t=t_{z}=0$ and $\mu=0$ are to be satisfied.
Figure S2: The evolution of the Hall conductivity $\sigma_{xy}$ (in unit of $e^{2}/h$) with respect to $B_{z}$ for different hopping amplitudes [(a,c)] and chemical potential [(b,d)]. We set $t=t_{z}$, $\mu=0$ in [(a,c)] and $t=t_{z}=0$ in [(b,d)]. The strength of spin-orbit coupling $\lambda_{\rm so}$ is set to 0 in [(a-b)] and $0.2$ in [(c-d)], respectively. Common parameters are $\lambda_{\rm M}=1$ and $\eta=1$.
@@
Intriguingly, if $t$ and $t_{z}$ remain zero, we observe that an approximate Hall plateau can emerge even for finite $\mu$, provided that the condition $B_{z}>\mu$ is satisfied. This phenomenon arises because the energy gaps in the $k_{z}$ planes (for $k_{z}\neq 0$ and $k_{z}\neq\pi$) increase as $B_{z}$. When $\mu$ lies within these energy gaps, the contributions from the $k_{z}$ planes become quantized. However, since this system is a semimetal, the energy gaps of the $k_{z}$ planes varies continuously as a function of $k_{z}$. A larger $B_{z}$ enhances the $k_{z}$-dependence of this function, making it more sharply varying. Nevertheless, planes sufficiently close to $k_{z}=0$ and $k_{z}=\pi$ always exhibit nonzero and nonquantized contributions. Consequently, the plateau is approximate as long as $\mu$ becomes finite.
@@
In real materials, $t$ and $t_{z}$ are generally finite, making the observation of the predicted Hall plateau less feasible. Nevertheless, a strong anomalous Hall effect induced by a weak field remains an intriguing and experimentally observable phenomenon. As a final remark, since the direction of the Zeeman field strongly affects both the symmetry and the band structure, the anomalous Hall effect exhibits a pronounced angular dependence. Specifically, finite components $\sigma_{xz}$ and $\sigma_{yz}$ may emerge when an in-plane magnetic field is applied and the combined symmetries $C_{2x}\mathcal{T}$ and $C_{2y}\mathcal{T}$ are broken. However, these components are not quantized because the two-dimensional momentum planes at fixed $k_{x}$ or $k_{y}$ correspond to trivial insulators when their spectra are gapped. This angular dependence provides an additional means to probe the band structure and the topology, complementing techniques such as angle-resolved photoemission spectroscopy and magneto-optical Kerr effects.
Figure S3: Canting of the local magnetic moments and anomalous Hall effects. (a-f) Schematic illustration of the orientations of the local magnetic moments in the absence [(a-c)] or presence [(d-f)] of a Zeeman field along the $z$ direction . The canting angle $\alpha$ ($\beta$) refers to the orientation deviation of the magnetic moments from $y$ ($x$) direction towards $z$ direction for magnetic atoms located in the yzyz (xzxz) planes of the unit cell. (g-f) Hall conductivity tensor $\sigma_{xy}$ as a function of $B_{z}/\lambda_{\rm M}$ for the full Hamiltonian in Eq. (S46). Without loss of generality, we set $\delta_{1x}=\delta_{1y}=0.1B_{z}$. In [(g)], the chemical potential is fixed at $\mu=0$, while in [(h)], spin-orbit coupling is turned off ($\lambda_{\rm so}=0$). Other parameters are $t=t_{z}=0$ and $\lambda_{\rm M}=\eta=1$.
IV IV. Zeeman field effects on the local magnetic moments
In the main article, we approximate the effect of an external magnetic field as introducing a Zeeman term for itinerant electrons. This approximation is well justified for weak magnetic fields, as in this regime, while the weak Zeeman term strongly influences the Berry curvature distribution, it has minimal impact on the magnetic background formed by local magnetic moments. In this section, we examine the canting of local magnetic moments (or spin flop) induced by the Zeeman coupling. This canting is accompanied by a weak ferromagnetic component. In principle, such canting does not occur in weak magnetic field because the exchange interaction in antiferromagnets is typically strong, and a comparable Zeeman energy (corresponding to a magnetic field on the order of 1-10 T) is required to reorient the spins. Nevertheless, for completeness, we analyze this effect here and demonstrate that, even for a weak spin canting, which itself already requires a large Zeeman field, the anomalous Hall effect discussed in Section III remains qualititively unchanged.
We begin with an energy scale analysis. Assume that the strength of magnetic field is less than 1T. The energy scale of the Zeeman term is of the order 0.1 meV (assuming the Landé factor $g=2$), which is much smaller than the typical energy scale of exchange interaction leading to high-temperature magnetic orders (e.g., 100 meV). This suggests that for materials whose magnetic order persists at relatively high temperatures (e.g., near room temperature), the influence of an external magnetic field on the magnetic order can generally be neglected. Nevertheless, how the local magnetic moment canting affects both the Hamiltonian and the AHE is an interesting question. To explore this effect, we consider a Zeeman field along the $z$ direction, enabling a direct comparison between results with and without including this effect.
A Zeeman field along the $z$-direction breaks symmetries including $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$ and $C_{4z}\mathcal{T}$ symmetries, while preserving $\mathcal{M}_{z}$, $\mathcal{M}_{x}\mathcal{T}$, and $\mathcal{M}_{y}\mathcal{T}$. The latter two symmetries indicate that the system remains invariant under the combination of a mirror reflection $\mathcal{M}_{x(y)}$ and time reversal $\mathcal{T}$. These conserved symmetries permit spin canting only along $z$, leading to a perturbation of the form: $\Delta\mathcal{H}(\boldsymbol{k})=\lambda_{\rm M}\sin k_{z}(\delta_{1x}\sin k_{x}+\delta_{1y}\sin k_{y})\sigma_{z}$, alongside a slight parameter renormalization of the spin-dependent hoppings in the $xz$ and $yz$ planes: $\lambda_{\rm M}\sin k_{z}(-\delta_{0x}\sin k_{x}\sigma_{x}+\delta_{0y}\sin k_{y}\sigma_{y})$. Physically, the parameters $\delta_{0x}$, $\delta_{1x}$} and $\delta_{0y}$, $\delta_{1y}$} should be functions of the canting angles $\beta$ and $\alpha$, respectively. We assume that the magnitude of each local magnetic moment is conserved, which imposes the normalization conditions: $\sqrt{\delta_{0x}^{2}+\delta_{1x}^{2}}=1$ and $\sqrt{\delta_{0y}^{2}+\delta_{1y}^{2}}=1$. A schematic of the orientations of these moments with and without canting is shown in Figs. S3 (a-f). Taking the perturbation into account, the full Hamiltonian becomes
$\mathcal{H}(\boldsymbol{k})$ $=$ $-\left[t(\cos k_{x}+\cos k_{y})+t_{z}\cos k_{z}\right]+\lambda_{\rm so}(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})$ (S46)
$+\lambda_{\rm M}\eta(\cos k_{x}-\cos k_{y})\sigma_{z}+\lambda_{\rm M}\sin k_{z}(-\delta_{0x}\sin k_{x}\sigma_{x}+\delta_{0y}\sin k_{y}\sigma_{y})$
$+\lambda_{\rm M}\sin k_{z}(\delta_{1x}\sin k_{x}+\delta_{1y}\sin k_{y})\sigma_{z}+B_{z}\sigma_{z}.$
Assuming a weak magnetic field, the canting angles $\alpha$, $\beta$ are expected to be small. Hence, without loss of generality, we assume that $\delta_{1x(y)}$ scales linearly with $B_{z}$. It is evident that the perturbation $\Delta\mathcal{H}(\boldsymbol{k})$ vanishes at momentum planes $k_{z}=\{0,\pi\}$, and does not introduce new band degeneracies. Thus, in the absence of SOC, i.e., $\lambda_{\rm so}=0$, each $k_{z}$-slice with $k_{z}\neq\{0,\pi\}$ remains a Chern insulator with the Chern number $\mathcal{C}(k_{z})=-{\rm sgn}(B_{z})$, yielding a quantized Hall conductivity $\sigma_{xy}=-\frac{e^{2}}{h}{\rm sgn}(B_{z})$. With finite SOC, the quantized Hall response persists as long as the $k_{z}$-positions of the Weyl nodes remain insensitive to $B_{z}$. The perturbation only slightly shifts the Weyl nodes within the $k_{z}$-planes given by: $\sin k_{x}(\lambda_{\rm so}-\lambda_{\rm M}\delta_{0x}\sin k_{z})=\sin k_{y}(\lambda_{\rm so}+\lambda_{\rm M}\delta_{0y}\sin k_{z})=0$. Numerical calculations [Figs. S3(g-h)] for this full Hamiltonian confirm that the Hall conductivity remains nearly unchanged (see the agreement between Fig. S2(b) and Fig. S3(h)). These results confirm that the canting effect to AHE can be safely neglected in weak magnetic fields.
V V. Connection to real-world materials
In this section, we discuss potential material realizations of the CNLs and magnetic Kramers Weyl nodes predicted by our model. In principle, realizing CNLs in real systems requires three mutually orthogonal mirror planes—a condition naturally satisfied in antiferromagnets with noncollinear magnetic moments belonging to orthorhombic, tetragonal, or cubic point groups. Such magnetic systems provide promising platforms for realizing the physics described in our work.
A particularly compelling candidate is the recently reported spin-split antiferromagnet $\rm MnTe_{2}$, identified by Zhu et al. [23]. Upon performing a coordinate transformation that redefines the in-plane momenta as $k_{1}=(k_{x}-k_{y})/\sqrt{2}$ and $k_{2}=(k_{x}+k_{y})/\sqrt{2}$, our model Hamiltonian maps onto a form that hosts band degeneracies consistent with those reported in $\rm MnTe_{2}$. The transformed Hamiltonian reads:
$\equiv$ $-\lambda_{\rm M}\left[\sqrt{2}\sin k_{z}\cos\frac{k_{x}}{\sqrt{2}}\sin\frac{k_{y}}{\sqrt{2}}\sigma_{x}+\sqrt{2}\sin k_{z}\sin\frac{k_{x}}{\sqrt{2}}\cos\frac{k_{y}}{\sqrt{2}}\sigma_{y}+2\eta\sin\frac{k_{x}}{\sqrt{2}}\sin\frac{k_{y}}{\sqrt{2}}\sigma_{z}\right],$
[…]
$H^{\Gamma}$ $=$ $\left[h_{1}+h_{3}(k_{x}^{2}+k_{y}^{2}+k_{z}^{2})\right]\sigma_{0}+h_{3}(k_{y}k_{z}\sigma_{x}+k_{z}{\color[rgb]{0,0,0}k_{x}}\sigma_{y}+k_{x}k_{y}\sigma_{z})+O(\boldsymbol{k}^{3}),$ (S48)
matches exactly with the low-energy expansion of our model when $\eta=1$:
$\mathcal{H}_{\rm CNL}^{\Gamma}$ $=$ $-\lambda_{\rm M}(k_{y}k_{z}\sigma_{x}+k_{x}k_{z}\sigma_{y}+k_{x}k_{y}\sigma_{z}),$ (S49)
demonstrating a one-to-one correspondence at low energies. Also, the absence of surface states on the $z$-normal surfaces, as reported in Ref. [23], is consistent with our theoretical prediction. Given these strong agreements, spin-ARPES measurements on the $x$- and $y$-normal surfaces of $\rm MnTe_{2}$ could be directly applied to detect the predicted surface states.
Turn to magnetic Kramers Weyl nodes. Since $\rm MnTe_{2}$ is centrosymmetric (belongs to point group $O_{h}$), spin-orbit coupled terms of the form $\lambda_{\rm so}\sum_{i=x,y,z}\sin k_{i}\sigma_{i}$ are forbidden, leading to the absence of magnetic Kramers Weyl nodes. However, symmetry-lowering perturbations—such as applied electric fields, uniaxial strain, interfacial engineering, or chemical doping—can induce such spin-orbit coupling while preserving $C_{4z}\mathcal{T}$ symmetry, thereby enabling magnetic Kramers Weyl nodes at symmetry-related TRIMs. Beyond $\rm MnTe_{2}$, chiral antiferromagnets such as $\rm Mn_{3}IrGe$ and $\rm YMnO_{3}$ have been predicted to host Weyl nodes at TRIMs by DFT calculations [103]. These materials also represent promising platforms for realizing the physics discussed here. Experimentally, magnetic Kramers Weyl nodes can manifest through long Fermi arcs (detectable via ARPES) and electromagnetic responses in linear and nonlinear regimes—including the anomalous Hall effect and circular photogalvanic effects. Notably, our model predicts quantized CPGE signals when the chemical potential is adjusted near the Weyl nodes.
Finally, we note that while exact CNL orthogonality arises in cubic, tetragonal, and orthorhombic systems, similar nodal degeneracies along high-symmetry lines may also arise in hexagonal and trigonal crystals, albeit with non-orthogonal configurations. Thus, the mechanisms uncovered in our study are applicable to a broad class of magnetic materials, spanning five of the seven crystal systems.